a man's wages increase from 625 naira to 7125 naira per month (1) find the increase percent (2) if the wages are taxed at 12%, find the increase in tax payment.
Answer by Theo(13342)

the wages increase from 625 to 7125.
the increase is 7125 - 625 = 6500.
the percent increase is 6500 / 625 = 10.4 * 100 = 1040 percent.

if the increase is taxed at 12 percent, then the increase in tax is.12 * 6500 = 780.

the original wages are 625, so the tax is .12 * 625 = 75
the increased wages are 7125, so the tax is .12 * 7125 = 855.
the increase in tax is 855 - 75 = 780.

your solutions are:
the increase in wages is 1040% of the original wages.
the increase in tax payment is 780.